OpenStudy (anonymous):

evaluate lim x → ∞ cos x 5 x

OpenStudy (john_es):

Do you mean \[\lim_{x\rightarrow \infty} \cos x\cdot 5x\]?

OpenStudy (anonymous):

yes

OpenStudy (dls):

<infinity

OpenStudy (john_es):

As cos x is bounded in [-1,1] you have \[constant\cdot\infty\] so it is infinty the result.